OpenStudy (anonymous):

There are seven black socks, eight blue socks, 10 white socks, and five patterned socks in the drawer. What is the probability of pulling a blue sock, replacing it, and then drawing a patterned sock out of the drawer?

OpenStudy (anonymous):

P(drawing blue) = 8/30 P( patterned) = 5/30 these probabilities are independent so we multiply P((blue then patterned) = 8/30 * 5 / 30 = 4/15 * 1/6 = 4/90 = 2/45

OpenStudy (anonymous):

required prob = (8/30) * (5/30) = 2/45

OpenStudy (anonymous):

Thanks guys :]

OpenStudy (anonymous):

yw